The University of California has two bonds outstanding. Both issues have the same credit rating, a face value of $1,000 and a coupon rate of 5%. Coupons are paid twice a year. Bond A matures in 1 year, while bond B matures in 30 years.

The market interest rate for similar bonds is 12%.

What is the price of bond A?

What is the price of bond B?

Now assume that yields increase to 15%. What is the price of bond A?

What is now the price of bond B?
Answer #1

Market rate -12%

Formula is =PV(rate,nper,pmt,fv)

Bond A:

=PV(12%/2,1*2,5%/2*1000,1000)

=935.83

Bond B:

=PV(12%/2,30*2,5%/2*1000,1000)

=434.35

Market rate -15%

Bond A:

=PV(12%/2,30*2,5%/2*1000,1000)

=434.35

Bond B:

=PV(15%/2,30*2,5%/2*1000,1000)

=342.03